OpenStudy (anonymous):

Which steps below show an equation that has NO solution? 5(x - 3) + 2 = -x + 47 5x - 15 + 2 = -x + 47 5x - 13 = -x + 47 5x = -x + 60 6x = 60 x = 10 6x - (6 + 4x) = 26 6x - 6 - 4x = 26 2x - 6 = 26 2x = 32 x = 16 3x + 9x - 5 = 2(6x + 1) - 7 3x + 9x - 5 = 12x + 2 - 7 12x - 5 = 12x - 5 -5 = -5 -4(1 - 6x) = 24x - 22 -4 + 24x = 24x - 22 24x = 24x -18 0 = 18 PLEASE I NEED HELP ON HOMEWORK

OpenStudy (loser66):

just for the last one. 0=18 ?? make sense??

OpenStudy (anonymous):

no its a full equation

OpenStudy (anonymous):

-4(1 - 6x) = 24x - 22 -4 + 24x = 24x - 22 24x = 24x -18 0 = 18

OpenStudy (anonymous):

i guess you are supposed to say the last step, where you get \(0=18\)

OpenStudy (anonymous):

oh okay thanks both of you :)

OpenStudy (anonymous):

btw i am not sure that is right i mean i am sure the equation has no solution, and i am sure you know it when you get \(0=18\) but the question is written in such a weird way, it is not clear what it wants either: the last step or: the whole thing

OpenStudy (anonymous):

okay :) thanks for answering :)